OpenStudy (anonymous):

solve for the solition set of the following expressions using completing square 1. x2+9x+18=0 2. 2x2+12x+4=0 3. x2+6x-8

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

1. 9/2 = 4.5 &4.5^2 = 20.25 so you need to add something onboth sides to 18 to get to 20.25 and then get a (x+4.5)^2= y 2. start by dividing the whole thing by 2